A chip used in a microcomputer contains 5.74 x 10-3g of silicon. Calculate the number of silicon atoms in the chip. [ Si = 28; Avogadros constant 6.02 x 1023]
1
Expert's answer
2021-01-18T01:57:52-0500

"m = 5.74×10^{-3}g\\
\textsf{molar mass} = 28g/mol\\
\textsf{number of moles} = \dfrac{5.74×10^{-34}}{28} = 2.05×10^{-4}\ moles"


1 mole of Si = 6.02 × 1023 atoms

2.05 × 10-4 moles of Si = x


x = "\dfrac{2.05×10^{-4}×6.02×10^{23}}1" = 1.234 × 1020 atoms.


Therefore, the chip has 1.234 × 1020 atoms.